OpenStudy (anonymous):

csc^2 x - 4 = 0, solve for x

OpenStudy (anonymous):

csc^2 x - 4 = 0 (csc x - 2)(csc x + 2) = 0 -> csc x - 2 = 0 or csc x + 2 = 0 -> csc x = 2 or csc x = -2 -> sin x = 1/2 or sin x = -1/2 I think you can solve it from there